OpenStudy (hopeless):

a 6.5 ft tall car standing next to an adult elephant casts a 33.2 ft shadow. If the adult elephant casts a shadow that is 51.5 ft long then how tall is it?

OpenStudy (anonymous):

take x as the height of the elephant. 6.5/33.2 = x/51.5 (6.5)(51.5) = 33.2x 33.2x = 334.75 x = 334.75/33.2 x = 10.0828 ft.